OpenStudy (anonymous):

What is the solution to the system of equations? 5x+4y=8 x-2y=10

OpenStudy (zpupster):

there are usually 2 ways used to solve these types of equation substitution or elimination since you can easily solve for x by adding 2y to both sides in 2nd equation do that that will give you x then substitute for x in first equation

OpenStudy (anonymous):

so... x=10? I think I did that wrong..

OpenStudy (zpupster):

5(2y+10) + 4y = 8 10y + 50 +4y =8 14y= 8-50 then resubstitute the y to get the x and check i have to run.